In LR series circuit L=20mh..R=5 ohm.. And E=6V.Calculate the time constant of the circuit and the current of the circuit after 5ms when the circuit is on
1
Expert's answer
2018-09-18T11:31:08-0400
The time constant of the circuit:
τ=L/R=(20·〖10〗^(-3))/5=4·〖10〗^(-3) s=4ms
The circuit after 5ms when the circuit is on:
I=I_0·(1-e^(- Rt/L) )=1.2·(1-e^(- (5·5·〖10〗^(-3))/(20·〖10〗^(-3) )) )=0.852 A
I_0=E/R=6/5=1.2 A